### v2

@@
[…]
Classically studied coherent structures only paint a partial picture of wall-bounded turbulence
[…]
Summary paragraph
For the last 140 years, the mechanisms of transport and dissipation of energy in a turbulent flow have not been completely understood due to the complexity of this phenomenon. The dissipation of energy due to turbulence is significative, and understanding turbulence physics is crucial for fighting the present climate emergency. Previous research has focused on analyzing the so-called coherent structures of the flow (Q events, streaks, and vortices), which are regions of high turbulence transport, high/low streamwise fluctuation, and rotation, respectively. However, the connection between these classically studied structures and the flow development is still uncertain. To fill this gap, here we show a data-driven methodology for objectively identifying high-importance regions in a turbulent flow. A deep-learning model is trained to predict a future state of a turbulent channel flow and the gradient-SHAP explainability algorithm is used to calculate the importance of each grid point for such a prediction. Finally, high-importance regions are computed using the SHAP data, analyzing and comparing their characteristics with those of the other coherent structures. The SHAP analysis provides an objective way to identify the regions of highest importance in the turbulent flow, which exhibit different levels of agreement with the classically studied structures.
@@
Although turbulent flows have been widely studied for over 140 years [1, 2, 3, 4], the nature of turbulence and its energy-transfer mechanisms remain unclear for most of the industrial and scientific flows [5, 6, 7, 8, 9]. A complete understanding of turbulence is crucial for the fight against the climate emergency [10]. In fact, 15% of the energy consumed worldwide is spent near the surface of vehicles due to turbulent effects [11].
Turbulent-flow dynamics are governed by the Navier–Stokes equations [12, 13], a set of partial differential equations, whose analytical solution for a general flow is probably impossible to obtain. The intricate behavior of these complex fluid flows is due to the multi-scale nature of turbulence [3, 14], which also imposes very high computational demands to simulate and study these flows [15]. The technique that numerically integrates the Navier–Stokes equations without any modeling is known as direct numerical simulation, DNS. Unfortunately, DNSs of practical situations are still decades away. It can be estimated that the random-access memory (RAM) needed to simulate a commercial jet is equivalent to that of a month of the internet.
As solving the governing equations is practically out of reach, many different techniques have appeared during these 140 years [16]. One of the most promising ideas is to understand, model, and finally control turbulence by studying the nonlinear interactions among the different coherent structures of the flow [17]. Historically, these coherent structures were first described experimentally as streamwise streaks [18] and Reynolds-stress events [19]. These two kinds of structures were defined in terms of the transport of kinetic energy and momentum. Later, in 1995 [20], a third class was proposed: vortices, which are defined by regions where vorticity was larger than shear. In all cases, the definitions of these structures are based on different physical effects. While the role of these structures in wall-bounded turbulence has been thoroughly documented, the literature lacks a completely objective assessment of the actual importance of these structures at different wall-normal locations.
@@
In a previous study, the intense Reynolds-stress structures were ranked depending on their impact on the evolution of a turbulent channel [27]. This work extends the previous results by calculating the importance of every single grid point in a turbulent channel flow in order to percolate that importance and define new importance-based structures by calculating the Shapley additive explanations (SHAP-based structures). For this purpose, an attribution method [28] based on the extension of the Shapley values [29] to an infinity-player game is required: the so-called gradient SHAP [30]. This algorithm implements the expected-gradients (EG) [31] methodology, a more computationally efficient version of the integrated gradients (IG) proposed by Sundararajan et al. [32].
@@
Figure 1: Diagram showing the methodology of the paper. The methodology is divided into three main stages. The first stage, represented by the black lines, comprises a U-net deep learning (DL) model for predicting the evolution of the flow. The three components of the simulated velocity fluctuation (green blocks, $u_{\rm{t}}$ for the streamwise, $v_{\rm{t}}$ for the wall-normal and $w_{\rm{t}}$ for the spanwise velocity) in an instant $t$ are evolved through the U-net to the next time step $t+1$ separated a viscous time $t^{+}=5$ (yellow blocks, $u_{\rm{t+1}}$, $v_{\rm{t+1}}$ and $w_{\rm{t+1}}$). In the second stage (blue flow connectors), the importance of each grid node of the field, for each velocity component, is calculated by applying the expected-gradients algorithm an extension of Shapley additive explanations (SHAP values) to an infinite-player game. For evaluating the importance of each input grid point in the prediction, the mean-squared error of the predicted field is used and the simulated fields in the next step (green blocks, $u_{\rm{t+1}}$, $v_{\rm{t+1}}$ and $w_{\rm{t+1}}$) is defined as the output. The result of this expectation is the so-called SHAP values, which are represented by the light-green blocks, and represent the importance of the streamwise ($\left(\phi_{u}\right)_{\rm{t}}$), wall-normal ($\left(\phi_{v}\right)_{\rm{t}}$) and spanwise ($\left(\phi_{w}\right)_{\rm{t}}$) velocity fluctuations. In the final stage (purple flow connectors), the previously mentioned SHAP values are used to segment the model through a percolation analysis. As a result, the flow field can be segmented into SHAP structures. Once these structures are defined, their physical characteristics are analyzed, and their correlation with other coherent structures is evaluated.
Definition of SHAP structures
@@
Based on the gradient-SHAP methodology we determine an importance score for each grid point of a turbulent channel with a friction Reynolds number $Re_{\tau}=u_{\tau}h/\nu=125$. The scores evaluate the influence of the grid points in the prediction of the state of the flow after a time interval $\Delta t^{+}=\Delta t\,u_{\tau}^{2}/\nu=5$, where $h$ is half height of the channel, $\nu$ is the fluid kinematic viscosity, $u_{\tau}=\sqrt{\tau_{\rm{w}}/\rho}$ is the friction velocity ($\tau_{\rm{w}}$ is the wall-shear stress and $\rho$ the fluid density) and $\Delta t$ the time interval between fields [16]. The friction Reynolds number sets the streamwise pressure gradient driving the flow, and thus, it is its main control parameter. The instantaneous velocity vector is represented by $\bm{U}$ with components $U$, $V$, and $W$ in their streamwise ($x$), wall-normal ($y$), and spanwise ($z$) components. The Reynolds decomposition is applied to the velocity, $U=\overline{U}+u$, where the statistically averaged quantities in $x$, $z$, and $t$ are denoted by an overbar, and the fluctuating quantities are denoted by lowercase letters. Additionally, the root mean square of the fluctuations is expressed with $(\cdot)^{\prime}$, and the magnitudes in viscous units are denoted by the superscript $+$. Additional details on the numerical approach and the simulation characteristics are provided in the Methods section.
@@
Figure 2: Instantaneous visualization of the various coherent structures in the channel flow. The figure shows four types of coherent structures in half of the channel colored by wall-normal distance (purple near the wall and yellow in the mid-plane of the channel): SHAP-based structures (top-left), Reynolds-stress structures or Q events (top-right), streaks (bottom-left) and vortices (bottom-right). In the figure, the wall is located at a wall-normal distance $y^{+}=0$, and the mid-plane of the channel is $y^{+}=125$. In addition, we use periodic boundary conditions on the planes $x^{+}=0$ and $x^{+}=1000\pi$, as well as on $z^{+}=0$ and $z^{+}=375\pi$. Note that the same flow field is used for the four panels.
@@
Once the gradient-SHAP methodology is applied to the problem, a vector field defining the importance of each region of the space in the evolution of the flow, namely the SHAP values $\phi$, is created. The SHAP vector provides the importance of the streamwise, wall-normal, and spanwise velocities, $\phi_{\rm{u}}$, $\phi_{\rm{v}}$, and $\phi_{\rm{w}}$, respectively. To establish a compact definition of what is considered a high-importance region, the combined effect of the three components of the vector is taken into account by comparing the absolute value of the vector against the square root of the sum of the square mean of each component as a function of the wall-normal distance:
@@
where $H$ is a percolation index set to a value of $2$ to maximize the total number of structures [34], as can be observed in the percolation block of Figure 1. A visual representation showing the SHAP-based structures, colored according to the wall-normal distance, is presented in the top-left image of Figure 2.
The definition of the SHAP-based structures is compared against some of the classical definitions of coherent structures widely used in the turbulence community. The first structures to compare with are the intense Reynolds-stress structures or Q events [35]. The structures are presented in the top-right panel of Figure 2 and are defined through the following expression:
@@
where $\alpha$ is the percolation parameter. The streamwise streaks are shown in the bottom-left panel of Figure 2. The low-velocity streaks result from the lift-up of the low-speed flow near the wall to the higher-velocity regions and they are complemented by the high-speed streaks, which redistribute the energy within the flow. Finally, the SHAP-based structures are also compared with the vortices defined by Chong et al. [20]. The definition of these structures is more complex and is presented in the Methods section. As a general idea, the vortices define those regions of the flow where rotation is larger than shear. The visualization of the vortices is provided in the bottom-right panel of Figure 2.
Analysis of the physical properties of SHAP-based structures
@@
Figure 3: Joint probability density function of the streamwise velocity fluctuation and the inner-scaled wall-normal distance for the different types of structures. The figure presents the SHAP-based structures (top-left), Reynolds-stress structures (top-right), streaks (bottom-left) and vortices (bottom-right). The histogram distribution contours are defined in the logarithmic scale, representing the most likely areas with yellow colors and the most unlikely ones with purple regions.
@@
Figure 3 shows that the SHAP-based structures exhibit several regions of high importance. The area of positive streamwise fluctuations in the viscous sublayer is in good agreement with the corresponding region of the Q events, which is associated with sweep structures or high-velocity structures moving towards the wall. Another area of importance is observed from $y^{+}\approx 30$ to the center of the channel, and this is also in good agreement with a region of high probability density of Q events, corresponding to low-velocity regions moving away from the wall, i.e. ejections. Note that sweeps and ejections have been widely studied in the literature as very important structures, and therefore the agreement with SHAP structures is a very important result. In fact, in a previous study, the ejections and the sweeps were demonstrated to exhibit a similar importance per unit of volume [27] and in the works of Jiménez [17] and Lozano-Durán et al. [35], the ejections and sweeps were highlighted as the most influential Q events. Furthermore, our results show a remarkable agreement between the wall-attached sweeps and the wall-detached ejections with the SHAP structures with positive and negative streamwise velocity fluctuations, respectively. This result is also consistent with the recent research of Osawa and Jiménez [21], which reinforces the idea of the sweeps being the most influential structure in the flow field as it transports the flow close to the wall, where the local shear is higher. At $y^{+}\simeq 15$ a strong agreement is observed between the SHAP structures and the streaks for both the positive and the negative streamwise fluctuations. The streaks are streamwise-elongated structures that play an important role in the turbulent flow by redistributing the energy between the low-velocity and the high-velocity regions. Therefore, the regions involving energy transfer between the wall-attached and wall-detached regions have been identified as SHAP structures with high importance. In Figure 3 a strong relationship between the high-importance regions and the regions of momentum transfer can be observed, as the Q events and streaks exhibit a stronger agreement with the SHAP structures, while the dissipative structures, such as the vortices, show lower agreement. Nevertheless, near the center of the channel ($y^{+}>50$) the region of high importance for zero streamwise velocity fluctuation $u\approx 0$ is explained by the presence of vortices and high-dissipation regions. However, although the SHAP structures exhibit high similarities with the previous structures in different regions of the channel, the distributions do not perfectly match any of them. Therefore, SHAP values detect complex nonlinear patterns that were not identified by the other definitions and approaches. For this reason, it is of high importance to analyze the percentage of agreement between the different structures as a function of the wall-normal distance to generate relationships between the properties of the coherent structures and their implications in the evolution of the flow.
@@
The coincidence of pairs of the various structures along the wall-normal direction is presented in Figure 4. This figure shows a roughly constant agreement between the SHAP structures and the intense Reynolds-stress structures (Qs). This co-occurrence represents around $60\%$ of the volume of the SHAP structures, and around $40\%$ of the volume of the intense Q events. This coexistence can be mainly observed in the red and yellow regions inside the black contours for the various wall-normal locations shown in Figure 4. Thus, a significant fraction of the SHAP structures coincides with Q events throughout the channel. However, there is a non-negligible percentage, 40% of the SHAP structures, which is not correlated with the Reynolds stress. This difference is due to two main reasons: on the one hand, the Q events are strongly related to the wall-normal velocity, while the SHAP structures generally give more importance to the streamwise component (the reader is referred to the Supplementary Material for a deeper analysis). On the other hand, the Q events do not take into account the spanwise fluctuations. Regarding the comparison between SHAP and streaks, for $10\lessapprox y^{+}\lessapprox 50$ the SHAP structures are contained inside the streaks, with almost $90\%$ agreement, as can be observed for $y^{+}\simeq 13$, where the black contours of the SHAP structures remain inside the large green streaks. Finally, the agreement between SHAP and vortices is around $20\%$ close to the channel center. Part of this coincidence is shared with the streaks and the Q events: as indicated by the orange, purple and light-blue colors inside the black contours. Although the regions containing only vortices are mostly outside the SHAP structures, a fraction of them are detected as highly influential for the evolution of the flow. This explains the agreement between the joint probability density functions of Figure 3 for SHAP and vortices when $u\simeq 0$ close to the channel center. Therefore, the SHAP regions represent a small fraction of the vortices, and, although the SHAP values are associated with some high rotation areas, this relationship is not direct.

### v3

@@
Abstract
For the last 140 years, the mechanisms of transport and dissipation of energy in a turbulent flow have not been completely understood. Previous research has focused on analyzing the so-called coherent structures, organized flow patterns characterized by their spatial coherence, lifespan and significant contribution to momentum and energy transfer. However, the connection between these structures and the flow development is still uncertain. Here, we show a data-driven methodology for objectively identifying high-importance regions. A deep-learning model is trained to predict a future state of the flow and the gradient-SHAP explainability algorithm is used to calculate the importance of each grid point. Finally, high-importance regions are computed using the SHAP data and are compared to the other coherent structures. The SHAP analysis provides an objective way to identify the regions of higher importance, which exhibit different levels of agreement with the classical structures without being completely related to any particular one.
@@ Abstract
Introduction
Although turbulent flows have been widely studied for over 140 years [1, 2, 3, 4], the nature of turbulence and its energy-transfer mechanisms remain unclear for most of the relevant industrial and scientific flows [5, 6, 7, 8, 9]. A complete understanding of turbulence may have worldwide implications, since 15% of the energy consumed worldwide is spent near the surface of vehicles due to turbulent effects [10].
Turbulent-flow dynamics are governed by the Navier–Stokes equations [11, 12], a set of non-linear partial differential equations, whose analytical solution for a general flow is probably impossible to obtain. The intricate behavior of these complex fluid flows is due to the multi-scale nature of turbulence [3, 13], which also imposes very high computational demands to simulate and study these flows [14]. The technique that numerically integrates the Navier–Stokes equations without any modeling is known as direct numerical simulation, DNS. Unfortunately, DNSs of practical situations are still decades away. It can be estimated that the random-access memory (RAM) needed to simulate a commercial jet is equivalent to that of a month of the internet.
As solving the governing equations is out of reach, many different techniques have appeared during these 140 years [15]. One of the most promising ideas is to understand, model, and finally control turbulence by studying the nonlinear interactions among the different coherent structures of the flow [16]. Historically, these coherent structures were first described experimentally as streamwise streaks [17] and Reynolds-stress events [18]. These two kinds of structures were defined in terms of the transport of kinetic energy and momentum. Later, in 1990 [19], a third class was proposed: vortices, which are defined by regions where vorticity was larger than shear. In all cases, the definitions of these structures are based on different physical effects. While the role of these structures in wall-bounded turbulence has been thoroughly documented, the literature lacks a completely objective assessment of the actual importance of these structures at different wall-normal locations.
@@ Introduction
Figure 1: Diagram showing the methodology of the paper. The methodology is divided into three main stages. The first stage, represented by the black lines, comprises a U-net deep learning (DL) model for predicting the evolution of the flow. The three components of the simulated velocity fluctuation (green blocks, $u_{\rm{t}}$ for the streamwise, $v_{\rm{t}}$ for the wall-normal and $w_{\rm{t}}$ for the spanwise velocity) in an instant $t$ are evolved through the U-net to the next time step $t+1$ separated a viscous time $t^{+}=5$ (yellow blocks, $u_{\rm{t+1}}$, $v_{\rm{t+1}}$ and $w_{\rm{t+1}}$). In the second stage (blue flow connectors), the importance of each grid node of the field, for each velocity component, is calculated by applying the expected-gradients algorithm an extension of Shapley additive explanations (SHAP values) to an infinite-player game. For evaluating the importance of each input grid point in the prediction, the mean-squared error of the predicted field is used and the simulated fields in the next step (green blocks, $u_{\rm{t+1}}$, $v_{\rm{t+1}}$ and $w_{\rm{t+1}}$) is defined as the output. The result of this expectation is the so-called SHAP values, which are represented by the light-green blocks, and represent the importance of the streamwise ($\left(\phi_{u}\right)_{\rm{t}}$), wall-normal ($\left(\phi_{v}\right)_{\rm{t}}$) and spanwise ($\left(\phi_{w}\right)_{\rm{t}}$) velocity fluctuations. In the final stage (purple flow connectors), the previously mentioned SHAP values are used to segment the model through a percolation analysis. As a result, the flow field can be segmented into SHAP structures. Once these structures are defined, their physical characteristics are analyzed, and their correlation with other coherent structures is evaluated for the different wall-normal distance $y^{+}$.
In a previous study, the intense Reynolds-stress structures were ranked depending on their impact on the evolution of a turbulent channel [26]. This work extends the previous results by calculating the importance of every single grid point in the flow in order to percolate that importance and define new importance-based structures by calculating the Shapley additive explanations (SHAP-based structures). For this purpose, an attribution method [27] based on the extension of the Shapley values [28] to an infinity-player game is required: the so-called gradient SHAP [29]. This algorithm implements the expected-gradients (EG) [30] methodology, a more computationally efficient version of the integrated gradients (IG) proposed by Sundararajan et al. [31].
@@
Results
@@
Based on the gradient-SHAP methodology we determine an importance score for each grid point of two turbulent channels, the first one with a friction Reynolds number $Re_{\tau}=u_{\tau}h/\nu=125$ and the second one with $Re_{\tau}=550$. These SHAP values identify the most influential points in the causal relationships detected by the deep-learning model. For this reason, the physical meaning of the SHAP values is conditioned by the output of the model. The SHAP values answer the question formulated mathematically for the deep-learning model. In the case presented in this work, the model calculates the mean-squared error of the reconstruction of the flow. Therefore, the SHAP values identify the most important regions of the flow for the evolution of the velocity fluctuation field. The scores evaluate the influence of the grid points in the prediction of the state of the flow after a time interval $\Delta t^{+}=\Delta t\,u_{\tau}^{2}/\nu=5$, where $h$ is half height of the channel, $\nu$ is the fluid kinematic viscosity, $u_{\tau}=\sqrt{\tau_{\rm{w}}/\rho}$ is the friction velocity ($\tau_{\rm{w}}$ is the wall-shear stress and $\rho$ the fluid density) and $\Delta t$ the time interval between fields [15]. The friction Reynolds number sets the streamwise pressure gradient driving the flow, and thus, it is its main control parameter. The instantaneous velocity vector is represented by $\bm{U}$ with components $U$, $V$, and $W$ in their streamwise ($x$), wall-normal ($y$), and spanwise ($z$) components. The Reynolds decomposition is applied to the velocity, $U=\overline{U}+u$, where the statistically averaged quantities in $x$, $z$, and $t$ are denoted by an overbar, and the fluctuating quantities are denoted by lowercase letters. Additionally, the root mean square of the fluctuations is expressed with $(\cdot)^{\prime}$, and the magnitudes in viscous units are denoted by the superscript ++. Further information on the numerical approach and the simulation characteristics are provided in the Methods section.
@@
Figure 2: Instantaneous visualization of the various coherent structures in the channel flow. The figure shows four types of coherent structures, SHAP-based structures, subfigures a) and b), Reynolds-stress structures or Q events, subfigures c) and d), streaks, subfigures e) and f) and vortices, subfigures g) and h), in half of the channel colored by wall-normal distance (purple near the wall and yellow in the mid-plane of the channel) for ${\rm Re}_{\tau}=125$, subfigures a), c), e) and g), and ${\rm Re}_{\tau}=550$, subfigures b), d), f) and h). In the figure, the wall is located at $y^{+}=0$ and the mid-plane of the channel is $y^{+}=125$ or $y^{+}=550$ depending on the case. In addition, we use periodic boundary conditions in $x$ and $z$ for both cases. Note that the same flow field is used for the four panels of each channel. Note that $x^{+}$, $y^{+}$ and $z^{+}$ define the streamwise, spanwise and wall-normal directions respectively.
@@
Once the gradient-SHAP methodology is applied to the problem, a vector field defining the importance of each region of the space in the evolution of the flow, namely the SHAP values $\phi$, is created. The SHAP values quantify the causal importance of each individual grid point in the evolution of the flow (refer to the supplementary material for a detailed discussion and a comparison with the causal system examples presented by Martínez-Sánchez et al. [33]). As previously mentioned, due to the evolution of the model, the SHAP values determine which grid points are the most influential for the evolution of the flow. Accordingly, the SHAP structures identify the regions of the flow that should be targeted to effectively control its evolution [34]. The SHAP vector provides the importance of the streamwise, wall-normal, and spanwise velocities, $\phi_{\rm{u}}$, $\phi_{\rm{v}}$, and $\phi_{\rm{w}}$, respectively. To establish a compact definition of what is considered a high-importance region, the combined effect of the three components of the vector is taken into account by comparing the absolute value of the vector against the square root of the sum of the square mean of each component as a function of the wall-normal distance:
@@
where $H$ is a percolation index set to a value of $2$ to maximize the total number of structures [35], as can be observed in the percolation block of Figure 1. A visual representation showing the SHAP-based structures, colored according to the wall-normal distance, is presented in subfigures a) and b) of Figure 2.
The definition of the SHAP-based structures is compared against some of the classical definitions of coherent structures widely used in the turbulence community. The first structures to compare with are the intense Reynolds-stress structures or Q events [36]. The structures are presented in subfigures c) and d) of Figure 2 and are defined as:
@@
where $\alpha$ is the percolation parameter. The streamwise streaks are shown in subfigures e) and f) of Figure 2. The low-velocity streaks result from the lift-up of the low-speed flow near the wall to the higher-velocity regions and they are complemented by the high-speed streaks, which redistribute the energy within the flow. These are regions of high turbulent kinetic energy. Finally, the SHAP-based structures are also compared with the vortices defined by Chong et al. [19]. The definition of these structures is more complex and is presented in the Methods section. In general, the vortices define those regions of the flow where rotation is larger than shear. Vortices appear close to the ejections [37] and contribute to the near-wall cycle to sustain turbulence involving the streaks. In fact, vortices are a result of the instabilities of the streaks and can be considered dissipative as they carry high levels of enstrophy. The visualization of the vortices is provided in subfigures g) and h) of Figure 2.
@@
Figure 3: Joint probability density function of the streamwise velocity fluctuation, $u^{+}$, and the inner-scaled wall-normal distance, $y^{+}$, for the different types of structures. The figure presents the SHAP-based structures, subfigures a) and b), Reynolds-stress structures, subfigures c) and d), streaks, subfigures e) and f) and vortices, subfigures g) and h) for ${\rm Re}_{\tau}=125$, subfigures a), c), e) and g), and ${\rm Re}_{\tau}=550$, subfigures b), d), f) and h).
@@
Figure 3 shows that the SHAP-based structures exhibit several regions of high importance for both friction Reynolds numbers. The area of positive streamwise fluctuations in the viscous sublayer is in good agreement with the corresponding region of the Q events, which is associated with sweep structures or high-velocity structures moving towards the wall. As presented in Figure 3, the high-importance high-velocity region close to the wall is similar for both friction Reynolds numbers, despite the higher separation of scales of ${\rm Re}_{\tau}=550$ with respect to ${\rm Re}_{\tau}=125$. In fact, our results show that the sweep events close to the wall constitute high-importance regions independently of the friction Reynolds number, as they transport the high-energy regions towards the wall, where the shear stress is stronger. Another area of importance is observed from $y^{+}\approx 30$ to the center of the channel, and this is also in good agreement with a region of high probability density of Q events, corresponding to low-velocity regions moving away from the wall, i.e. ejections. Although the high-importance regions are similar for SHAP structures and Q events at ${\rm Re}_{\tau}=125$, larger differences are observed for ${\rm Re}_{\tau}=550$. While the SHAP structures only detect a region of high importance for low-velocity events, the Q events increase the probability of sweeps. Thus, the SHAP structures can identify the Q events with higher impact for the evolution of the flow, or in other words, the most causal Q events. Note that sweeps and ejections have been widely studied in the literature as very important structures, and therefore the agreement, as well as the differences, with SHAP structures are a very important result. In fact, in a previous study, the ejections and the sweeps were demonstrated to exhibit a similar importance per unit of volume [26] and in the works of Jiménez [16] and Lozano-Durán et al. [36], the ejections and sweeps were highlighted as the most influential Q events. Furthermore, our results show a remarkable agreement between the wall-attached sweeps and the wall-detached ejections with the SHAP structures with positive and negative streamwise velocity fluctuations, respectively. This result is also consistent with the recent research of Osawa and Jiménez [20], which reinforces the idea of the sweeps being the most influential structure in the flow field as it transports the flow close to the wall, where the local shear is higher. At $y^{+}\simeq 15$ a strong agreement is observed between the SHAP structures and the streaks for both the positive and the negative streamwise fluctuations. The streaks are streamwise-elongated structures that play an important role in the turbulent flow by redistributing the energy between the low-velocity and the high-velocity regions. Therefore, the regions involving energy transfer between the wall-attached and wall-detached regions have been identified as SHAP structures with high importance for both friction Reynolds numbers. In Figure 3 a strong relationship between the high-importance regions and the regions of momentum transfer can be observed, as the Q events and streaks exhibit a stronger agreement with the SHAP structures, while the dissipative structures, such as the vortices, show lower agreement. Nevertheless, for $y^{+}>50$ the region of high importance for zero streamwise velocity fluctuation $u\approx 0$ is explained by the presence of vortices and high-dissipation regions. However, although the SHAP structures exhibit high similarities with the previous structures in different regions of the channel, the distributions do not perfectly match any of them. Therefore, SHAP values detect complex nonlinear patterns that were not identified by the other definitions and approaches, as could be clearly observed with the Q events far from the wall for ${\rm Re}_{\tau}=550$. For this reason, it is of high importance to analyze the percentage of agreement between the different structures as a function of the wall-normal distance to generate relationships between the properties of the coherent structures and their implications in the evolution of the flow.
Figure 4: Coincidence between the various types of structures. Percentage of coincidence of pairs of the following structures: SHAP, Q events, streaks and vortices, relative to the volume of each type of the pair for ${\rm Re}_{\tau}=125$, a), and ${\rm Re}_{\tau}=550$, b), along wall-normal distance $y^{+}$.
@@
The coincidence of pairs of the various structures along the wall-normal direction is presented in Figure 4. This figure shows a roughly constant agreement between the SHAP structures and the intense Reynolds-stress structures (Qs). This co-occurrence represents around $60\%$ of the volume of the SHAP structures, and around $40\%$ of the volume of the intense Q events for ${\rm Re}_{\tau}=125$ and around $50\%$ and $30\%$ respectively for ${\rm Re}_{\tau}=550$. This coexistence can be mainly observed in the blue and green regions inside the black contours for the various wall-normal locations shown in Figure 5. Thus, a significant fraction of the SHAP structures coincides with Q events throughout the channel for both frictions Reynolds numbers, but specially for ${\rm Re}_{\tau}=125$. However, there is a non-negligible percentage, 40% and 50% for each friction Reynolds number respectively of the SHAP structures, which is not correlated with the Reynolds stress. This difference is due to two main reasons: on the one hand, the Q events are strongly related to the wall-normal velocity, while the SHAP structures generally give more importance to the streamwise component (the reader is referred to the Supplementary Material for a deeper analysis). On the other hand, the Q events do not take into account the spanwise fluctuations. Regarding the comparison between SHAP and streaks, for $10\lessapprox y^{+}\lessapprox 50$ the SHAP structures are contained inside the streaks, with almost $90\%$ agreement in both cases, as can be observed for $y^{+}\simeq 13$, where the black contours of the SHAP structures remain inside the large orange or green streaks. Finally, the agreement between SHAP and vortices is lower and remains below $20\%$ in both cases. Moreover, part of this coincidence is shared with the streaks and the Q events: as indicated by the purple, brown and gray colors inside the black contours. Although the regions containing only vortices are mostly outside the SHAP structures, a fraction of them are detected as highly influential for the evolution of the flow. This explains the agreement between the joint probability density functions of Figure 3 for SHAP and vortices when $u\simeq 0$ close to the channel center. Therefore, the SHAP regions represent a small fraction of the vortices, and, although the SHAP values are associated with some high rotation areas, this relationship is not direct.